{
  "single_attempt_probability": 0.01,
  "assumptions": [
    "binary outcome",
    "independence",
    "fixed p"
  ],
  "formula": "P(at least one success)=1-(1-p)^n",
  "values": [
    {
      "attempts": 1,
      "at_least_one": 0.01,
      "all_failed": 0.99
    },
    {
      "attempts": 10,
      "at_least_one": 0.09561792499119552,
      "all_failed": 0.9043820750088045
    },
    {
      "attempts": 50,
      "at_least_one": 0.39499393286246337,
      "all_failed": 0.6050060671375366
    },
    {
      "attempts": 100,
      "at_least_one": 0.6339676587267705,
      "all_failed": 0.3660323412732295
    },
    {
      "attempts": 230,
      "at_least_one": 0.9008951844811252,
      "all_failed": 0.09910481551887486
    },
    {
      "attempts": 459,
      "at_least_one": 0.9900790257989593,
      "all_failed": 0.009920974201040622
    },
    {
      "attempts": 1000,
      "at_least_one": 0.9999568287525893,
      "all_failed": 4.3171247410658195e-05
    }
  ],
  "plot_n_range": [
    0,
    500
  ]
}
